paul spend his income of 100$ per week on two goods : pizza ( which cost $8each ) and milk( which cost $1 per gallon) at his current level of consumption , the marginal utility from the last pizza consumed is 32 and the marginal utility from the last gallon of milk is 4. is colin maximizing his utility? why or why not? if not , what should paul do to achieve a higher level of utility ?

1
Expert's answer
2021-11-25T10:18:40-0500

Thus, the utility maximization rule from the point of view of marginal utility theory is that:

The consumer must spend his income in such a way that the ratio of marginal utility to price is the same for all goods, while the income must be spent completely:

"\frac{MUx}{Px} = \frac{MUy}{Py}"

where MU is the marginal utility of goods X and Y, and P is their price.

The MU/P ratio shows the value of marginal utility per 1 ruble spent.

"\frac{32}{8} = \frac{4}{1}"

4=4

Therefore, maximizes its usefulness